OpenStudy (anonymous):

Find the vertex of the parabola for the equation: 3x2 - 6x - 5 A. (1, -4) B. (1, -8) C. (-1, -4) D. (-1, -8)

OpenStudy (callisto):

3x^2 - 6x - 5 =3(x^2-2x) -5 =3(x^2-2x+1-1)-5 = 3(x-1)^2 -3-5 = 3(x-1)^2 -8 For quadratic equation having the form y=a(x-h)^2 +k (h,k) are the coordinates of the vertex. Compare 3(x-1)^2 -8 with a(x-h)^2 +k Can you get the vertex now?

OpenStudy (anonymous):

Yes Thank You

OpenStudy (callisto):

welcome!